OpenStudy (tileer139):

What is the slope of the line passing through the points (2, 5) and (0, −4)?

OpenStudy (tileer139):

@Mehek14

OpenStudy (tileer139):

so 9\2

OpenStudy (tileer139):

medal and fan for help

pooja195 (pooja195):

9/2 is right. Welcome to OpenStudy! :-)

OpenStudy (tileer139):

thx

OpenStudy (tileer139):

but how

pooja195 (pooja195):

\(\dfrac{5-(-4)}{2-0}\) = ? 2 -'s turn into a postivie \[\huge~\rm~\bf~ \frac{ 5+4=9 }{ 2-0=2 }\]

OpenStudy (liv1234):

There is a formula for a slope passing through a line. m=y2-y1/x2-x1

OpenStudy (tileer139):

i thought the 2 was over the 9

pooja195 (pooja195):

Mehek made a mistake (2, 5) and (0, −4) Should've been -4-5/0-2 This is the slope formula \[\LARGE\color{red}{ \frac{y_2-y_1}{x_2-x_1}}=slope\] \[\huge~\rm~\bf~ \frac{ -4-5 }{ 0-2 }=?\]

OpenStudy (tileer139):

so -9/-2

pooja195 (pooja195):

Yes however slopes can be simplifed -/-=+ So... \[\huge~\rm~\bf~ \color{purple}{\frac{ -9 }{ -2 }\rightarrow \frac{ 9 }{ 2}}\]

pooja195 (pooja195):

We still got the same answer lololol

OpenStudy (tileer139):

can u help with one more

pooja195 (pooja195):

yes :-)

OpenStudy (tileer139):

What is the slope of the line passing through the points (–1, 3) and (4, –7)? A. B. C. D.

pooja195 (pooja195):

Try this one by yourself tell me what you get :)

OpenStudy (tileer139):

2 3/4 -4/3 -2 are the answers

OpenStudy (tileer139):

i dont know it i was in the hs when it was taought at school

OpenStudy (tileer139):

my test is timed

pooja195 (pooja195):

Ok, So we have the slope formula \[ \huge~\rm~\bf\color{blue}{ \frac{y_2-y_1}{x_2-x_1}=slope}\] We are given the following coordinates: (–1, 3) and (4, –7) x1=-1 x2=4 y1=3 y2=-7 Now we plug what we have into the formula \[ \huge~\rm~\bf\color{purple}{ \frac{-7-3}{4--1}=slope}\] two negatives make a + so.. \[ \huge~\rm~\bf\color{red}{ \frac{-7-3}{4+1}=slope}\] Now solve, what would our slope be?

OpenStudy (tileer139):

-10/5

OpenStudy (tileer139):

is wat i got

pooja195 (pooja195):

Good slopes can be simplifed so.. -10 divided by 5 =?

OpenStudy (tileer139):

so -2

pooja195 (pooja195):

Yes! ^.^ Perfect :)

OpenStudy (tileer139):

thanks

pooja195 (pooja195):

yw ^_^